# A Public Dataset For the ZKsync Rollup

### v1

@@ Abstract.
Despite blockchain data being publicly available, practical challenges and high costs often hinder its effective use by researchers, thus limiting data-driven research and exploration in the blockchain space. This is especially true when it comes to Layer 2 (L2) ecosystems, and ZKsync, in particular.
To address these issues, we have curated a dataset from 1 year of activity extracted from a ZKsync Era archive node and made it freely available to external parties. In this paper, we provide details on this dataset and how it was created, showcase a few example analyses that can be performed with it, and discuss some future research directions. We also publish and share the code used in our analysis on GitHub to promote reproducibility and to support further research.
1. Introduction
@@
The blockchain ecosystem is built on the principles of decentralization and transparency. However, there remains a significant challenge for end users and non-technical researchers: blockchain data is not easily accessible. This challenge is hindering the widespread adoption of blockchain technology, which should be more easily accessible.
@@
Currently, blockchain data can be obtained by deploying an archive node in the case of chains based on Ethereum Virtual Machine (EVM) (ZKs, 2024d; Offchain Labs, 2024; Optimism Foundation, 2024; bnbchain.org, 2024), or a full node for Bitcoin (bitcoin.org, 2024). However, this requires a machine with high bandwidth and high-speed storage to keep the blockchain fully synced. In other words, individual deployment of archive nodes is not a practical solution. Therefore, it would be advantageous if we could offer an alternative that improves this situation. Additionally, users can gather the data through Remote Procedure Calls (RPCs)providers. However, gathering data this way can be challenging for some non-technical users and researchers and is usually costly. Alternatively, users can depend on external data sources such as Etherscan (Etherscan, 2024), Arbiscan (Arbitrum, 2024), Dune (Dune Analytics, 2024), ZettaBlock (ZettaBlock Inc., 2024), Nansen (Nan, 2024), and similar platforms. Although they may provide an interesting solution, they may prove to be costly in the long run, and do not meet the requirements of certain end-users, particularly those engaged in research that relies on easily and cheaply accessible data.
@@
We believe that anyone requiring blockchain data should have a straightforward and hassle-free way to access them, without worrying about infrastructure or hardware. This data has significant value for research purposes, such as alerting (Messias et al., 2021b), Airdrop designs (Messias et al., 2023b), analyzing Maximal-Extractable Value (MEV) (Messias et al., 2023a; Gogol et al., 2023, 2024c), Automated Market Maker (AMM) (Gogol et al., 2024a), and potentially improving the adoption of particular chains such as ZKsync and other Layer-2 blockchains (L2s)rollups within the research scope (Messias et al., 2024a; Gogol et al., 2024d). In addition, it can contribute to the growth of these ecosystems as more researchers become involved in the field.
@@
In a recent talk, a company called Paradigm emphasized that to empower the blockchain community with the necessary data analysis capabilities, we must make blockchain data more accessible (Ret, 2024). By ensuring that these data are available affordably, quickly, and with minimal effort, we can effectively address the challenge of data unavailability in our ecosystem. They introduced a novel EVMblockchain node developed in Rust, known as Reth that can be applied to EVM-compatible blockchains. Paradigm provided an endpoint to their archive node allowing anyone to make requests to their node. Similar initiatives have been provided by Matter Labs and other companies interested in making blockchain data more accessible (ZKs, 2024d; Offchain Labs, 2024; Optimism Foundation, 2024), Nevertheless, users would need to build high-efficiency code to gather all the data they need, which could invalidate this process for some users due to network delays or lack of coding skills.
@@
1.1. Why ZKsync Era Data?
Launched in March 2023, ZKsync Era is a L2scaling solution for the Ethereum blockchain that utilizes Zero-Knowledge Proof (ZKP)for efficient transaction processing. ZKsync Era is also an EVM-based rollup, ensuring compatibility with existing Ethereum smart contracts. As of July 2024, ZKsync Era ranks among the top five ZKPchains with a Total Value Locked (TVL)estimated at 1.23 billion USD (L2Beat, 2024). This scaling solution improves Ethereum by processing transactions in batches using ZKP, which helps maintain low transaction fees and encourages user participation.
Currently, rollups are a key strategy to scale the Ethereum ecosystem (Buterin, 2020). In fact, in the last few years, these L2chains have become an important part of the ecosystem by pushing new innovations and absorbing a significant portion of user activity. At the same time, there are still many unexplored questions when it comes to these blockchains, and existing research is fairly limited.
Given the growing role of L2chains in the Ethereum ecosystem, we believe that there is an opportunity for researchers to conduct research and expand our understanding of ZKsync. By making our ZKsync Era data easily available to external groups, we hope to advance ZKsync-related research and knowledge about L2chains, more generally.
1.2. Contributions
@@
1.3. Paper Organization
This paper is organized as follows. Section 2 details our dataset and its data schema, together with the necessary background to understand the context of blockchain data in general. Section 3 provides examples of analyses that can be achieved using this dataset such as regarding transaction fees and gas usage, events derived from transaction logs, token swaps, and more. Furthermore, Section 4 discusses open problems and future directions for which this dataset could be useful. Finally, we present the conclusion of our work in Section 5.
@@ 1.3. Paper Organization
02-data-exploration-fees.ipynb It provides the analyses regarding gas usage and transaction fees for ZKsync used in Section 3.1. We use two main sources of data: blocks and transaction receipts.
@@
04-data-exploration-swaps.ipynb It analyzes the swap events on ZKsync described in Section 3.3. We use one main source of data: transaction logs but also load blocks data to extract timestamps information.
Table 1. Description of the code used for analysis in this paper. These notebook files are available on our GitHub repository (ZKs, 2024a) in the directory ./zksync-data-dump/notebooks/ and provide examples of how to interact and process our ZKsync public dataset.
2. Data Schema and Processing
@@
Table 2. Description of our 1-year ZKsync dataset.
We gathered our dataset from our ZKsync Era archive node as raw data. This data consists of all the information regarding blocks, transactions, receipts, and logs. Then, to allow anyone to use the dataset, we conducted a pre-processing step. This step consists of formatting the data in a parquet format that can be easily accessible through well-known libraries available in Python, for example, Pandas (Pandas, 2024) and Polars (Polars, 2024b). Due to the high volume of data (i.e., around 200 GB), we focus on using Polars for better processing, memory management, and Lazy evaluation (Johnsson, 1984; Polars, 2024a). Next, we discuss each of the different data types separately.
2.1. Blocks
Blocks are sequential units of data within a blockchain, each identified by a unique hash. They contain a list of transactions, metadata such as timestamps, and the hash of the previous block (parentHash), which links them in a chain back to the genesis block (block number 0) as shown in Figure 1. This chain of blocks forms the blockchain. Blocks ensure transaction security, network consensus, and efficient data storage and processing within blockchain networks. We list the attributes of the blocks presented in our dataset in Table 3 in the Appendix. In our dataset, blocks should be sorted according to the attribute block number to maintain the correct order of the blocks on the blockchain.
Figure 1. Illustration of blocks connected to each other forming the blockchain.
2.2. Transactions
Transactions are digital interactions that involve the transfer of assets, the recording of data, or the execution of smart contracts between parties on blockchains like those based on EVM. Each transaction is initiated by a user, authenticated through cryptographic signatures, and sent to a decentralized network of nodes for validation. Once verified, transactions are grouped into blocks and added to the blockchain via a consensus mechanism, ensuring that they are secure, immutable, transparent, and free of intermediaries, forming the core of the blockchain system. We list the attributes of the transactions presented in our dataset in Table 4 in the Appendix.
@@ 2.2. Transactions
2.3. Transactions Receipts
Besides the raw transaction data described above, our dataset also contains a table of transaction receipts.
Transaction receipts provide a comprehensive summary of the outcome and effects of a transaction once it is processed and included in a block. They include the transaction hash, block number, and block hash to identify and verify the transaction, along with the sender (from) and recipient (to) addresses.
The receipts also detail the cumulative gas used by the transaction and all preceding transactions in the block, the actual gas used by the specific transaction, and the final gas price paid by the user. By multiplying the actual gas used by the gas price, we have the actual transaction fee the user paid.
Additional information includes logs for event logging (discussed next), the transaction status (success or failure), and the effective gas price paid. These receipts are crucial for users and developers to understand, audit, and interact with transactions and smart contracts on the blockchain. For example, they provide essential elements for the analysis, monitoring, and verification of transaction fees spent by users. We list the attributes of the transactions presented in our dataset in Table 5 in the Appendix.
@@ 2.4. Transactions Logs
These logs are emitted using the emit keyword within the smart contract code and play a pivotal role in monitoring activities, triggering actions within decentralized applications, and enabling event-driven programming. For instance, Decentralized Exchanges (DEXs)emit events upon trade executions, enabling user interfaces to update displays with current trade information.
These logs are stored in transaction receipts, offering a gas-efficient method to capture transient event data without permanently altering the blockchain’s state. Among the vast array of data accessible on EVM-based blockchains, transaction logs stand out as crucial sources of information for researchers, developers, and users. They facilitate analyses of various token transfer patterns and support blockchain analysis research, the focus of our contributions.
@@ 2.4.1. Topics Attributes
2.4.2. Hashing and Signatures
topics0 corresponds to the hashed function signature using keccak256 (Bertoni et al., 2009). This signature consists of the function name followed by its parameter types. For example, the signature of a typical Transfer event is Transfer(address,address,uint256). After hashing it with keccak256, the result becomes 0xddf2 $\cdots$ b3ef, which is the topics0. Below is a Python code snippet demonstrating how to verify if a given signature matches topics0:
@@ 2.4.2. Hashing and Signatures
2.4.3. Event Mapping
We provide a mapping of the most frequently invoked events within the ZKsync dataset in our GitHub repository (ZKs, 2024a) under ./src/utils.py#events_dict. This mapping facilitates the parsing of the majority of events in our dataset. The mapping is structured as a dictionary where the topics0 hex value serves as the key, and the corresponding value is a dictionary containing the parsed event name and its function signature. For instance, the Transfer event is represented as follows within the map: [^1] We shortened topics0 to 0xddf252ad $\cdots$ f523b3ef for better visualization in the paper.
1events_dict[’0xddf252ad ... f523b3ef’] = {
2 ’name’: ’Transfer’,
3 ’signature’: ’Transfer(address,address,uint256)’}
2.5. L2 to L1 Logs
L2to L1logs are messages emitted by the ZKsync L2network and transmitted to the Ethereum L1network. They are essential for maintaining communication between the two layers, ensuring the security and integrity of transactions and data transfers. In ZKsync, the L1smart contract verifies these communications by checking the messages alongside the ZKP. The only provable part of the communication from L2to L1is the native L2to L1logs emitted by the Virtual Machine (VM). These logs can be generated using the to_l1 opcode (ZKs, 2024c). We refer the reader to the ZKsync documentation for more details (ZKs, 2024b). We list the attributes of these logs in Table 7 in §A in the appendix.
3. Example Analyses
@@
This section should serve as a starting point for other researchers wishing to use this dataset for their investigations. The code for generating this analysis is discussed in Table 1 and can be accessed on our GitHub repository (ZKs, 2024a).
3.1. Gas Usage and Transaction Fees
@@
Figure 2 shows the daily transactions executed in ZKsync Era over the period analyzed. The network processed an average of $905194$ transactions daily, with a significant spike in December 2023 that $5362921$ transactions in a single day. This spike was due to a boom in inscriptions, which became very popular for memecoin traders around this time.  Messias et al. (2024a) did a comprehensive review of this phenomenon across various rollups. Since the spike stabilized, we have seen a slight increase in daily transactions, which are now hovering around 1.2 million transactions per day.
Figure 2. Total transactions executed per day.
Now, looking at gas usage, Figure 3 illustrates the total number of gas units used each day in the ZKsync Era. On average, the network has processed $567111979658$ gas units per day in the last year. However, gas usage has been fairly volatile. Particularly, it has experienced three significant spikes of more than 1.5 trillion gas units per day, which we detail next.
May 2023.  30% of all gas used during the month of May 2023 can be attributed to the zkApes airdrop. The address receiving the most gas units was the SyncSwap router contract, which is used to execute swaps in the largest DEXin ZKsync Era. During this time, we see more than 10% of all swaps trading ZAT, which does not occur when we look at all swaps during the year under analysis. The second address receiving the most gas units was the contract used to claim the zkApes token (ZAT) in the context of the airdrop.
December 2023.  This was caused by the inscriptions boom we discussed before.
@@
Figure 3. Total gas units used per day.
Figure 4. Average gas units used per transaction.
We can further contextualize gas utilization by looking at Figure 4, which shows the average number of gas units used per transaction. We see that in May 2023, transactions were using on average a significantly higher number of gas units (reaching $7265323$ units per transaction) compared to the average over the 1-year period ($784149$ units).
Finally, we can look at the transaction fees. Figure 5 shows the average transaction fees over time. In other words, this is the cost that users have paid, on average, to submit one unit of gas to ZKsync Era. Transaction fees were fairly stable during 2023, with an average of 0.25 gwei per unit of gas.
However, in 2024, a key upgrade led to a significant decrease in ZKsync transaction fees. Concretely, the implementation of the new prover, Boojum (Matter Labs, 2024), marked a significantly reduced hardware requirement to run a Zero-Knowledge (ZK)prover and thus allowed a reduction of transaction fees to 0.1 gwei per unit of gas.
Then, in March 2024, the Dencun upgrade was implemented and deployed on the Ethereum mainnet. This upgrade, brought in EIP-4844, introduced a new type of transaction that can store “blobs” of data in the beacon node for 14 days (Vitalik Buterin and Dankrad Feist and Diederik Loerakker and George Kadianakis and Matt Garnett and Mofi Taiwo and Ansgar Dietrichs, 2022). Blobs have their independent fee model and submitting data as blobs is much cheaper than the previously used call-data (Omkar Godbole, 2024). ZKsync Era was one of the first rollups using blobs to publish the data related to its state changes, thus further reducing transaction fees to 0.025 gwei per unit of gas.
Figure 5. Average transaction fee in gwei per unit of gas by day.
3.2. Events and Contract Deployments
Events are an important source of additional data about activity on EVMchains. They correspond to data emitted by smart contracts and stored on-chain. In this subsection, we analyze this data and report on some specific events. Recall that events can be obtained from the logs data in the ZKsync dataset as discussed in Section 2.
Figure 6 shows the top 15 event types with the most emitted events during the period under analysis. Different contracts may emit the same event type, so what distinguishes them is the event’s function signature, and thus its respective hash.
There is a significant overrepresentation of the top 4 event types, with Transfer events being by far the largest type (70.9% of all events are Transfer events). This is not unexpected as this event is emitted every time an ERC-20 token is transferred between two addresses. This occurs in both simple token transfers, but also as part of other common contract operations such as swaps in DEXs.

### v2

# Maria Inês Silva 1144    Johnnatan Messias 22    Benjamin Livshits 33

@@ Abstract
Despite blockchain data being publicly available, practical challenges and high costs often hinder its effective use by researchers, thus limiting data-driven research and exploration in the blockchain space. This is especially true when it comes to Layer-2 (L2) ecosystems, and ZKsync, in particular. To address these issues, we have curated a dataset from 1 year of activity extracted from a ZKsync Era archive node and made it freely available to external parties. We provide details on this dataset and how it was created, showcase a few example analyses that can be performed with it, and discuss some future research directions.
1 Introduction
@@
The blockchain ecosystem is built on the principles of decentralization and transparency. However, a significant challenge remains for end users and non-technical researchers: blockchain data is not easily accessible. This challenge hinders the widespread adoption of blockchain technology, which should be more easily accessible.
@@
Currently, blockchain data can be obtained by deploying an archive node in the case of chains based on Ethereum Virtual Machine (EVM) [8, 12, 30, 33], or a full node for Bitcoin [11]. However, this requires high bandwidth and a high-speed storage machine to keep the blockchain fully synced. In other words, individual deployment of archive nodes is not a practical solution. Additionally, users can gather the data through Remote Procedure Calls (RPCs)providers. However, collecting data this way can be challenging for some non-technical users and researchers and is usually costly. Alternatively, users can depend on external data sources such as Etherscan, Arbiscan, Dune, ZettaBlock, Nansen, and similar platforms. Although they may provide an interesting solution, they may prove costly in the long run and do not meet the requirements of specific end-users, particularly those engaged in research that relies on quickly and cheaply accessible data.
@@
We believe that anyone requiring blockchain data should have a straightforward and hassle-free way to access it without worrying about infrastructure or hardware. This data has significant value for research purposes, such as alerting and measurement analysis [20, 25], Airdrop designs and analyzes [15, 29], analyzing Maximal-Extractable Value (MEV) [18, 19, 27, 39, 42, 45], Automated Market Maker (AMM) [16, 40], and potentially improving the adoption of particular chains such as ZKsync and other Layer-2 blockchains (L2s)rollups within the research scope [26]. In addition, it can contribute to the growth of these ecosystems as more researchers become involved in the field.
@@
In a recent talk, a company named Paradigm emphasized that to empower the blockchain community with the necessary data analysis capabilities, we must make blockchain data more accessible [6]. By ensuring that these data are available affordably, quickly, and with minimal effort, we can effectively address the challenge of data unavailability in our ecosystem. They introduced a novel EVMblockchain node developed in Rust, known as Reth, that can be applied to EVM-compatible blockchains. Paradigm provided an endpoint to their archive node, allowing anyone to make requests to their node. Similar initiatives have been provided by Matter Labs and other companies interested in making blockchain data more accessible [8, 30, 33]. Nevertheless, users would need to build high-efficiency code to gather all the necessary data, which could invalidate this process for some users due to network delays or lack of coding skills.
@@
1.1 Why ZKsync Era Data?
Launched in March 2023, ZKsync Era is a L2scaling solution for the Ethereum blockchain that utilizes Zero-Knowledge Proof (ZKP)for efficient transaction processing. ZKsync Era is also an EVM-based rollup, ensuring compatibility with existing Ethereum smart contracts. In July 2024, ZKsync Era ranked among the top five ZKPchains with a Total Value Locked (TVL)estimated at 1.23 billion USD [23]. This scaling solution improves Ethereum by processing transactions in batches using ZKP, which helps maintain low transaction fees and encourages user participation.
In this sense, rollups are a key strategy to scale the Ethereum ecosystem [13]. In fact, these L2chains have become an important part of the ecosystem in the last few years by pushing new innovations and absorbing a significant portion of user activity. At the same time, there are still many unexplored questions when it comes to these L2 blockchains, and existing research on them is fairly limited.
Given the growing role of L2chains in the Ethereum ecosystem, we believe there is an opportunity for researchers to conduct research and expand our understanding of these L2 blockchains, such as ZKsync. By making a 1-year ZKsync Era data easily available to external groups, we hope to advance ZKsync-related research and knowledge about L2chains, more generally.
1.2 Contributions
@@
1.3 Paper Organization
This paper is organized as follows. Section 2 details our dataset and its data schema, together with the necessary background to understand the context of blockchain data in general. Section 3 provides examples of analyses that can be achieved using this dataset, such as regarding transaction fees and gas usage, events derived from transaction logs, token swaps, and more. Furthermore, Section 4 discusses open problems and future directions for which this dataset could be useful. Finally, we present the conclusion of our work in Section 5.
Table 1: Description of the code used for analysis in this paper. These notebook files are available in our GitHub repository [4] in the directory ./zksync-data-dump/notebooks/ and show how to interact and process our ZKsync dataset.
@@ 1.3 Paper Organization
02-data-exploration-fees.ipynb It analyses gas usage and transaction fees for ZKsync used in Section 3.1. We use two main sources of data: blocks and transaction receipts.
@@
04-data-exploration-swaps.ipynb It analyzes the swap events on ZKsync described in Section 3.3. We use one main source of data: transaction logs but also load blocks data to extract timestamps.
2 Data Schema and Processing
@@
We gathered our dataset from a ZKsync Era archive node as raw data. This data consists of all the information regarding blocks, transactions, receipts, and logs. Then, we conducted a pre-processing step to allow anyone to use the dataset. This step consists of formatting the data in a parquet format that can be easily accessible through well-known libraries available in Python, for example, Pandas [34] and Polars [38]. Due to the high volume of data (i.e., around 200 GB), we focus on using Polars for better processing, memory management, and Lazy evaluation [22, 37], allowing straightforward data processing on a local laptop. Next, we discuss each of the different data types separately.
2.1 Blocks
Blocks are sequential units of data within a blockchain, each identified by a unique hash. They contain a list of transactions, metadata such as timestamps, and the hash of the previous block (parentHash), which links them in a chain back to the genesis block (block number 0) as shown in Figure 1. This chain of blocks forms the blockchain. Blocks ensure transaction security, network consensus, and efficient data storage and processing within blockchain networks. In our dataset, blocks should be sorted according to the attribute block number to maintain the correct order of the blocks on the blockchain.
Figure 1: Illustration of blocks connected to each other forming the blockchain.
2.2 Transactions
Transactions are digital interactions that involve the transfer of assets, the recording of data, or the execution of smart contracts between parties on blockchains like those based on EVM. Each transaction is initiated by a user, authenticated through cryptographic signatures, and sent to a decentralized network of nodes for validation. Once verified, transactions are grouped into blocks and added to the blockchain via a consensus mechanism, ensuring that they are secure, immutable, transparent, and free of intermediaries, forming the core of the blockchain system.
@@ 2.2 Transactions
2.3 Transactions Receipts
Besides the transaction data, our dataset also contains a table of transaction receipts. These transaction receipts provide a comprehensive summary of the outcome and effects of a transaction once it is processed and included in a block. They include the transaction hash, block number, and block hash to identify and verify the transaction, along with the sender (from) and recipient (to) addresses. The receipts also detail the cumulative gas used by the transaction and all preceding transactions in the block, the actual gas used by the specific transaction, and the final gas price paid by the user. By multiplying the actual gas used by the gas price, we have the actual transaction fee the user paid.
Additional information includes logs for event logging (discussed next), the transaction status (success or failure), and the effective gas price paid. These receipts are crucial for users and developers to understand, audit, and interact with transactions and smart contracts on the blockchain. For example, they provide essential elements for analyzing, monitoring, and verifying transaction fees spent by users.
@@ 2.4 Transactions Logs
These logs are emitted using the emit keyword within the smart contract code and play a pivotal role in monitoring activities, triggering actions within decentralized applications, and enabling event-driven programming. For instance, Decentralized Exchanges (DEXs)emit events upon trade executions, enabling user interfaces to update displays with current trade information. These logs are stored in transaction receipts, offering a gas-efficient method to capture transient event data without permanently altering the blockchain’s state. Among the vast array of data accessible on EVM-based blockchains, transaction logs stand out as crucial sources of information for researchers, developers, and users. They facilitate analyses of various token transfer patterns and support blockchain analysis research, which is the focus of our contributions.
@@
2.4.2 Hashing and Signatures.
topics0 corresponds to the hashed function signature using keccak256 [10]. This signature consists of the function name followed by its parameter types. For example, the signature of a typical Transfer event is Transfer(address,address,uint256). After hashing it with keccak256, the result becomes 0xddf2 $\cdots$ b3ef, [^1] We shortened topics0 to 0xddf252ad $\cdots$ f523b3ef for better visualization in the paper. which is the topics0. Below is a Python code snippet demonstrating how to verify if a given signature matches topics0:
@@
2.4.3 Event Mapping.
We provide a mapping of the top 90 most frequently invoked events within the ZKsync dataset in our GitHub repository under ./src/utils.py#events_dict. This mapping facilitates the parsing of the majority of events in our dataset. The mapping is structured as a dictionary where the topics0 hex value serves as the key, and the corresponding value is a dictionary containing the parsed event name and its function signature. For instance, the Transfer event is represented as follows within the map:
1events_dict["0xddf252ad1be2c89b69c2b068fc378daa952ba7f163c4a11628f55a4df523b3ef"] = {
2 "name": "Transfer",
3 "signature": "Transfer(address,address,uint256)"}
2.5 L2 to L1 Logs
L2to L1logs are messages emitted by the ZKsync L2network and transmitted to the Ethereum L1network. They are essential for maintaining communication between the two layers, ensuring the security and integrity of transactions and data transfers. In ZKsync, the L1smart contract verifies these communications by checking the messages alongside the ZKP. The only provable part of the communication from L2to L1is the native L2to L1logs emitted by the Virtual Machine (VM). These logs can be generated using the to_l1 opcode [7]. We refer the reader to the ZKsync documentation for more details [5].
3 Example Analyses
@@
This section should serve as a starting point for other researchers wishing to use this dataset for their research. The code for generating this analysis is discussed in Table 1 and will be available in a GitHub repository.
3.1 Gas Usage and Transaction Fees
@@
Figure 2 shows the daily transactions executed in ZKsync Era over the period analyzed. The network processed an average of $905194$ transactions daily, with a significant spike in December 2023 that accounts for $5362921$ transactions in a single day. This spike was due to a boom in inscriptions, which became very popular for memecoin traders around this time. Messias et al. [26] did a comprehensive review of this phenomenon across various rollups. Since the spike stabilized, we have seen a slight increase in daily transactions, which are now hovering around 1.2 million transactions per day.
Figure 2: Total transactions executed per day.
Considering gas usage, Figure 3(a) illustrates the total number of gas units used each day in the ZKsync Era. On average, the network has processed $567111979658$ gas units per day in the last year. However, gas usage has been relatively volatile. Particularly, it has experienced three significant spikes of more than 1.5 trillion gas units per day, which we detail next.
May 2023.  30% of all gas used during the month of May 2023 can be attributed to the zkApes airdrop. The address receiving the most gas units was the SyncSwap router contract, which is used to execute swaps in the largest DEXin ZKsync Era. During this time, we see more than 10%
December 2023.  This was caused by the inscriptions boom we discussed before [26].
@@
(a) Total gas units.
(b) Average gas unit.
Figure 3: Comparison of gas unit usage over time: (a) Total gas units used per day; and (b) Average gas units used per transaction.
We can further contextualize gas utilization by looking at Figure 3(b), which shows the average number of gas units used per transaction. We see that in May 2023, transactions were using on average a significantly higher number of gas units (reaching $7265323$ units per transaction) compared to the average over the 1-year period ($784149$ units).
Finally, we can look at the transaction fees. Figure 4 shows the average transaction fees over time. In other words, this is the cost that users have paid, on average, to submit one unit of gas to ZKsync Era. Transaction fees were fairly stable during 2023, with an average of 0.25 gwei per gas unit.
However, in late 2023, a key upgrade led to a significant decrease in ZKsync transaction fees. Concretely, the implementation of the new prover, Boojum [24], marked a significantly reduced hardware requirement to run a Zero-Knowledge (ZK)prover and thus allowed a reduction of transaction fees to 0.1 gwei per unit of gas.
Then, in March 2024, the Dencun upgrade was implemented and deployed on the Ethereum mainnet. This upgrade, brought in EIP-4844, introduced a new type of transaction that can store “blobs” of data in the beacon node for 14 days [44]. Blobs have their independent fee model and submitting data as blobs is much cheaper than the previously used call-data [32]. ZKsync Era was one of the first rollups using blobs to publish the data related to its state changes, thus further reducing transaction fees to 0.025 gwei per unit of gas.
Figure 4: Average transaction fee in gwei per unit of gas by day.
3.2 Events and Contract Deployments
Events are an important source of additional data about activity on EVMchains. They correspond to data emitted by smart contracts and stored on-chain. Next, we analyze this data and report on some specific events. Recall that events can be obtained from the logs data in the ZKsync dataset as discussed in Section 2.
Figure 5 shows the top 15 event types with the most emitted events during the period under analysis. Different contracts may emit the same event type, so what distinguishes them is the event’s function signature and, thus, its respective hash.
There is a significant overrepresentation of the top 4 event types, with Transfer events being by far the most significant type (70.9% of all events are Transfer events). This is not unexpected as this event is emitted every time an ERC-20 token is transferred between two addresses in ZKsync Era. This occurs in simple token transfers and other standard contract operations such as swaps in DEXs.